OpenStudy (anonymous):

I have a problem involving slope and y intercept. The problem relates to the following: A 5 mile cab ride costs $7.20 and a 9 mile ride costs $11.60 I calculated the slope, which is 1.1 but how do I find the y intercept? When I graph the coordinates, the line doesn't cross the y axis?

OpenStudy (anonymous):

Ok you know the slope, and you can think of the data as points on a graph (5,7.20) (9,11.60) So put it into point slope form y-y1=m(x-x1) where (x1,y1) is a point you choose to use and m is the slope. Then simplify

OpenStudy (anonymous):

That part I get but I need to write a linear equation for y = mx + b with the b part being the y intercept.

OpenStudy (anonymous):

Yes, thats what you will get when you plug the values in and simplify the equation i wrote.

OpenStudy (anonymous):

Okay. Thanks for your help!